OpenStudy (anonymous):

Which are measurements of the sides of a right triangle? Select one of the options below as your answer: A. 202, 198, 40 B. 28, 26, 12 C. 17, 14, 6 D. all of the above

OpenStudy (ranga):

Square each number. See if the biggest square is the sum of the two smaller squares. If yes, then it is a right triangle. Otherwise, no.

OpenStudy (ranga):

202^2 = 40,804 198^2 = 39204 40^2 = 1600 198^2 + 40^2 = 39204 + 1600 = 40,804 which equals 202^2 and therefore this is a right triangle. Same procedure for the other two choices.

OpenStudy (ranga):

B) 28, 26, 12 Square each number: 28 x 28 = 784 26 x 26 = 676 12 x 12 = 144 Add the two small squares: 676 + 144 = 820. This does NOT match the biggest square 784. Therefore, this is NOT a right triangle.

OpenStudy (ranga):

You can try C. Same procedure.